ACCURATE POTENTIAL ENERGY SURFACE, ROVIBRATIONAL ENERGY LEVELS, AND TRANSI-TIONS OF AMMONIA C_{3v} ISOTOPOLOGUES: ^{14}NH_{3}, ^{15}NH_{3}, ^{14}ND_{3} and ^{14}NT_{3}

\underline{XINCHUAN HUANG}, MS 245-6, NASA Ames Research Center, Moffett Field, CA, 94035; DAVID W. SCHWENKE, MS T27B-1, NASA Ames Research Center, Moffett Field, CA, 94035; TIMOTHY J. LEE, MS 245-1, NASA Ames Research Center, Moffett Field, CA, 94035.

A further refined, global potential energy surface (PES) is computed for the C_{3v} symmetry isotopologues of ammonia, including ^{14}NH_{3}, ^{15}NH_{3}, ^{14}ND_{3} and ^{14}NT_{3}. The refinement procedure was similar to that used in our previously reported PES^{a}, but now extends to higher J energy levels and other isotopologues. Both the diagonal Born-Oppenheimer correction and the non-adiabatic correction were included. J = 0 − 6 rovibrational energy levels and transition frequencies of ^{14}NH_{3} computed on this PES are in excellent agreement with HITRAN data. Statistics on nearly 4100 transitions and more than 1000 energy levels demonstrate the accuracy achieved by the state-of-the-art ”Best Theory + Experiment” strategy. Most transition frequencies are of ±0.01−0.02 cm^{−1} accuracy. Similar accuracy has been found on^{15}NH_{3} J = 0−3 rovibrational energy levels. Several transitions and energy levels in HITRAN have been identified as unreliable or suspicious, and some have been re-assigned. For ^{14}ND_{3} and ^{14}NT_{3}, J = 0 − 3 calculations have been performed. Agreement for pure rotationinversion transitions is nearly perfect, with more reliable energy levels presented. On the other hand, our J = 0 results suggest a re-analysis on the ^{14}ND_{3} ν_{1} band origin is needed. Finally, we will discuss possible future refinements leading to an even better final PES for Ammonia.

aX. Huang, D.W. Schwenke, and T.J. Lee, J. Chem. Phys. \underline{129}, 214304 (2008).